# Achronal localization and representation of the causal logic from conserved current, application to massive scalar boson

### v2

@@ 4 Flux passing through a maximal achronal set
The result (4) essentially is an extension of [12, Proposition 37] to spacelike Cauchy surfaces. Note that [12, Proposition 37] is equally valid for a $C^{1}$ vector field in place of a smooth one. General spacetime theories study Cauchy surfaces. These are the sets which meet every inextendible timelike smooth curve exactly once [20, Chapter 14, Definition 28]. By [20, Chapter 14, Lemma 29] they meet even all inextendible causal smooth curves. According to [7, (9) Remark, (36) Theorem] a spacelike Cauchy surface is just a causal base, i.e., it is a spacelike [^2] A set is called spacelike according to sec. 2. set which intersects every causal line. It equals the graph of the corresponding $1$-Lipschitz function.
(9) Lemma. Let $\tau:\mathbb{R}^{3}\to\mathbb{R}$ and $S:=\{(\tau(x),x):x\in\mathbb{R}^{3}\}$. Then $S$ is a spacelike Cauchy surface if $|\tau(x)-\tau(y)|<|x-y|$ for $x\neq y$ and $\limsup_{|x|\to\infty}|\tau(x)|/|x|<1$ or if a fortiori $\tau$ is $L$-Lipschitz with $L<1$.
@@
(10) Theorem. Let the $C^{1}$ vector field $\mathfrak{j}$ on $\mathbb{R}^{4}$ be bounded and satisfy the continuity equation (4.1). Suppose that it is zero or causal future-directed, i.e.
@@
       (b) The claim is that $\mathfrak{h}$ is locally Lipschitz. By [12, Proposition 37], $\Phi:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}^{4}$, $\Phi(s,x):=\gamma_{x}(s)$ is a diffeomorphism. Write $\Phi=:(\varphi,\phi)$.
       From the implicit function theorem [16] applied to the function $f:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}$, $f:=\varphi-\tau\circ\phi$, it follows that $\sigma:\mathbb{R}^{3}\to\mathbb{R}$ is continuous and hence that $\mathfrak{h}$ is continuous. Indeed, since $f(s,x)=0$ is uniquely solved by $s=\sigma(x)$ for every $x\in\mathbb{R}^{3}$ it remains to verify that $f(\cdot,x)$ is injective for every $x\in\mathbb{R}^{3}$. Let $f(s,x)=f(s^{\prime},x)$. This implies $|\varphi(s,x)-\varphi(s^{\prime},x)|\leq|\phi(s,x)-\phi(s^{\prime},x)|$. Since $\gamma_{x}$ is timelike, $s=s^{\prime}$ follows.
       Now we show that $\mathfrak{h}$ is even locally Lipschitz. Put $\mathfrak{G}:=(\varphi,\operatorname{id}_{\mathbb{R}^{3}})$. As $\partial_{s}\varphi(s,x)=v_{0}(\Phi(s,x))>0$ one infers that $\mathfrak{G}$ is an diffeomorphisms. Check $g(h_{0}(x),x)=h(x)$ for $g:=\phi\circ\mathfrak{G}^{-1}$.
       Fix $x,y\in\mathbb{R}^{3}$. Define $\beta:[0,1]\to\mathbb{R}^{3}$, $\beta(u):=g(\mathfrak{z}(u))$ with $\mathfrak{z}(u):=\big{(}(h_{0}(y),y)-(h_{0}(x),x)\big{)}u+(h_{0}(x),x)$. Check $\beta(1)=h(y)$, $\beta(0)=h(x)$ and remind $\mathfrak{h}(\mathbb{R}^{3})=S$.
@@
       Now $\beta^{\prime}(u)=D_{time}g(\mathfrak{z}(u))\big{(}(h_{0}(y)-h_{0}(x)\big{)}+D_{space}g(\mathfrak{z}(u))\big{(}y-x\big{)}$, whence $|h(y)-h(x)|\leq||D_{time}g(\mathfrak{z}(u_{0}))||\,|h(y)-h(x)|+|D_{space}g(\mathfrak{z}(u_{0}))|(y-x)$.
@@
       Let $K\subset\mathbb{R}^{3}$ be a compact ball. The image of $K$ under the continuous map $(h_{0},\operatorname{id}_{\mathbb{R}^{3}})$ is contained in a compact ball $D\subset\mathbb{R}^{4}$. In conclusion, $|h_{0}(y)-h_{0}(x)|\leq C_{D}|y-x|$ for all $x,y\in K$.
       (c) Let $0<R<\infty$. Put $A_{R}:=\{x\in\mathbb{R}^{3}:|x|=R\}$ and $B_{R}:=\{x\in\mathbb{R}^{3}:|x|<R\}$. Suppose that $\sigma(x)>0$ for $|x|<R$. Then $E_{R}:=\bigcup_{x\in B_{R}}\{\gamma_{x}(s):0<s<\sigma(x)\}\subset\mathbb{R}^{4}$ is open with $\partial E_{R}=M\cup M_{0}$, where $M:=\{0\}\times B_{R}\cup L_{R}\cup\mathfrak{h}(B_{R})$ with $L_{R}:=\bigcup_{x\in A_{R}}\{\gamma_{x}(s):0<s<\sigma(x)\}$, and $M_{0}:=\{0\}\times A_{R}\cup\mathfrak{h}(A_{R})$.
       One verifies that (3) applies to $n=4$, $E=E_{R}$ showing $\mathcal{H}^{3}(\partial E_{R})<\infty$ by (d), (e), and $\mathcal{M}^{3}(M_{0})=0$ by (f), (3).
       (d) Let $C\subset\mathbb{R}^{3}$ be bounded Borel. Then $\mathcal{H}^{3}(\{0\}\times C)<\infty$, $\mathcal{H}^{3}(\operatorname{graph}\tau|_{C})<\infty$ by [17, Theorem 8.1].
@@
       Recall the diffeomorphism $\Phi:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}^{4}$. Hence $f:=\mathbb{R}\times]-R/2,R/2[\times]-R/2,R/2[\to\mathbb{R}^{4}$, $f(s,x_{1},x_{2}):=\Phi\big{(}s;x_{1},x_{2},\sqrt{R^{2}-x_{1}^{2}-x_{2}^{2}}\,\big{)}$ is injective, $C^{1}$ whence Lipschitz on compact sets, and $f([-S,S]\times[-R/3,R/3]\times[-R/3,R/3])=D^{\prime}$. The claim holds by [17, Theorem 8.1].
       (f) Let $C\subset\mathbb{R}^{3}$ be a compact Lebesgue null set. Then $\mathcal{H}^{3}(\{0\}\times C)=0$ and $\mathcal{H}^{3}(\mathfrak{h}(C))=0$ hold by [17, Theorem 8.1]. Indeed, the former is obvious, the latter holds true as $\mathfrak{h}$ is Lipschitz on $C$ being compact.
       (g) Now let $\eta\in C_{c}^{1}(\mathbb{R}^{4})$ with $0\leq\eta\leq 1$ and $\eta|_{\overline{E}_{R}}=1$. Put $\phi:=\eta\mathfrak{v}$. Then (3) yields $0=\int_{M}\mathfrak{v}\,\nu_{E_{R}}\operatorname{d}\mathcal{H}^{3}$ as $\mathfrak{v}$ satisfies the continuity equation. Moreover the integration over $L_{R}$ yields 00 as the integrand is 00. So using [17, Theorem 9.1] it follows $\int_{\varpi(\mathfrak{h}(B_{R}))}\big{(}v_{0}(\tau(x),x)-v(\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}x^{3}=\int_{B_{R}}v_{0}(0,x)\operatorname{d}x^{3}$.
       (h) Finally, as explained in [12, Proposition 37], the condition $\sigma(x)>0$ for $|x|<R$ in (c) can be removed and the limit $R\to\infty$ carried out. One ends up with (4.2).∎
Under the assumptions made on the vector field $\mathfrak{j}$, the result in (4) states that the future-directed flux passing through a spacelike Cauchy surface is the same for all these surfaces. Under the additional assumption (4)(b) on $\mathfrak{j}$ the result in (4) is extended to all maximal achronal sets.
(11) Corollary. Let the $C^{1}$ vector field $\mathfrak{j}$ be bounded and satisfy the continuity equation (4.1). Suppose that for all $\mathfrak{x}\in\mathbb{R}^{4}$
@@
    $j_{0}(\mathfrak{x})\geq|j(\mathfrak{x})|$,
@@
    $\gamma\,j_{0}(\mathfrak{x})+k(x)\geq\,|j(\mathfrak{x})|$ for some nonnegative integrable bounded $C^{1}$ function $k:\mathbb{R}^{3}\to\mathbb{R}$ and some $10<\gamma<$ 1.
Then for every maximal achronal set $\Lambda$ being the graph of the corresponding $1$-Lipschitz function $\tau:\mathbb{R}^{3}\to\mathbb{R}$
@@
holds. The integrands are nonnegative and the integrals may be infinite.
Proof. Due to (a) and $|\nabla\tau|\leq 1$ the integrands are nonnegative. Define the vector field $\mathfrak{l}=(l_{0},l)$ by $l_{0}(\mathfrak{x}):=\gamma j_{0}(\frac{1}{\gamma}x_{0},x)+k(x)$, $l(\mathfrak{x}):=j(\frac{1}{\gamma}x_{0},x)$.
       One easily checks $(\partial_{i}l_{i})(\mathfrak{x})=(\partial_{i}j_{i})(\frac{1}{\gamma}x_{0},x)$ for $i=0,1,2,3$. Hence $\mathfrak{l}$ satisfies the continuity equation.
       Obviously the function $\rho:=\gamma\tau$ is $\gamma$-Lipschitz. As $\gamma<1$, by (4) the graph of $\rho$ is a spacelike Cauchy surface.
       So (4) applies to $\mathfrak{l}$ and $\rho$, whence
$\int\big{(}l_{0}(\rho(x),x)-l(\rho(x),x)\nabla\rho(x)\big{)}\operatorname{d}^{3}x=\int l_{0}(0,x)\operatorname{d}^{3}x\>.$
Replace $l_{0},l$ and $\rho$ by their definitions and multiply both sides by $\frac{1}{\gamma}$. One obtains (4.3) up to the finite positive summand $\frac{1}{\gamma}\int k(x)\operatorname{d}^{3}x$ on both sides.∎
Obviously one has
(12) Lemma. The assumption (4)(b) is satisfied, if (4)(a) and $j_{0}(x_{0},x)\leq\frac{1}{1-\gamma}k(x)$ hold.
@@ 5 Covariant achronal localization out of covariant conserved current
       There exist AL with the localization operators $T(\Delta)$ being orthogonal projections [7, (22) Theorem]. In this case the localization operators commute. However, a quantum mechanical system localized by a projection valued AL necessarily does not have a semi-bounded energy operator. This no-go result following from Hegerfeldt ’s well-known theorem regards a first type of Einstein causality requirement whose modern generalized reformulation [5, 7] we shall present below.
       The notion of localization of the above definition seems appropriate to describe measurement processes where a quantum system is absorbed by the apparatus and no further localization measurments can performed on it. This is because, in case of subsequent measurements a second type of Einstein’s causality requirement essentially regarding the no-signaling condition has to be considered. To comply with this requirement, under some popular assumptions about the post measurement state, $T(\Delta)$ and $T(\Delta^{\prime})$ should commute if $\Delta$ and $\Delta^{\prime}$ cannot be joined by causal curves. This requirement cannot be fulfilled as a consequence of Malament’s theorem and its modern re-formulations [^3] A quick review on these issues related to various facets of Einstein’s causality and localization appears in the introduction of [18]..
@@
In [12], condition CC was proved true in a special case where the sets $\Delta$ belong to smooth Cauchy surfaces and in [7], the final result has been established that an AL satisfies CC in full. As argued in [7], CC even necessitate achronal localization. In fact spacelike localization is not sufficient since CC induces the localization in achronal hyperplanes. This fact is reported in [7] and studied in detail in [8]. Hence we consider the study and the explicit construction of the achronal localization of the massive scalar boson to be of utmost relevance.
       It is the very principle of causality which let one think of the probability of localization as a conserved quantity reigned by an associated density current $\mathfrak{J}$. Indeed, by (5) a covariant AL can be constructed by means of a covariant conserved current assuming that on the Euclidean space its zeroth component describes the density of the probability of localization, namely explicitly $J_{0}(\phi;0,x)\geq 0$, $x\in\mathbb{R}^{3}$, and
@@
(14) Definition. Let $\mathcal{D}$ be a $W$-invariant dense subspace of $\mathcal{H}$. Let $\mathfrak{J}=(J_{0},J)$ be a map from $\mathcal{D}\times\mathbb{R}^{4}$ to $\mathbb{R}^{4}$ such that $\mathfrak{J}(\phi,\cdot)$ is a bounded $C^{1}$ vector field. (i) $\mathfrak{J}$ is conserved if the latter satisfies the continuity equation. (ii) $\mathfrak{J}$ is (Poincaré) covariant if $\mathfrak{J}\big{(}W(g)\phi,\mathfrak{x}\big{)}=A\cdot\mathfrak{J}\big{(}\phi,g^{-1}\cdot\mathfrak{x}\big{)}$, $g=(\mathfrak{a},A)\in\tilde{\mathcal{P}}$ holds.
@@
(19) Theorem. Let $\mathfrak{J}$ be a conserved covariant current. Let $\phi\in\mathcal{D}$. Let $J_{0}$ satisfy
@@
    $J_{0}(\phi,\mathfrak{x})\leq C_{\phi}(1+x^{2})^{-2}$ for some $0<C_{\phi}<\infty$
Then there is a covariant AL $T$ satisfying for every achronal Borel set $\Delta$
@@
where $\tau:\varpi(\Delta)\to\mathbb{R}$ with $\operatorname{graph}\tau=\Delta$.
@@
Let $\phi\in\mathcal{D}$. By (a) and (5), $J_{0}(\phi,\mathfrak{x})\geq|J(\phi,\mathfrak{x})|$. Hence by (b) and (4), $\frac{1}{2}J_{0}(\phi,\mathfrak{x})+\frac{1}{2}C_{\phi}(1+x^{2})^{-2}\geq|J(\phi,\mathfrak{x})|$.

### v4

@@ 4 Flux passing through a maximal achronal set
The result (4) essentially is an extension of [13, Proposition 37] to spacelike Cauchy surfaces in Minkowski spacetime. Note that [13, Proposition 37] is equally valid for a $C^{1}$ vector field in place of a smooth one. General spacetime theories study Cauchy surfaces. These are the sets which meet every inextendible timelike smooth curve exactly once [25, Chapter 14, Definition 28]. By [25, Chapter 14, Lemma 29] they even meet all inextendible causal smooth curves. According to [8, (9) Remark, (36) Theorem] a spacelike Cauchy surface in Minkowski spacetime is just a causal base, i.e., it is a spacelike [^2] A set is called spacelike according to sec. 1. set which intersects every causal line. It equals the graph of the corresponding $1$-Lipschitz function.
       The basic idea of the proof of the crucial technical result (4) is the application of the Divergence Theorem to a distorted cylindrical surface in $\mathbb{R}^{4}$ (which is $\partial E_{R}$ in the proof of (4)) with a flat bottom base, an achronal top base, and the side constituted by integral curves of the divergence-free $C^{1}$ vector field. As the flux through the side is zero, the fluxes through the bases must be equal. This were the desired result. The crux is that the versions of the Divergence Theorem known to us do not apply to this kind of surface, which obviously is not polyhedral, nor $C^{1}$ as the top base need not be $C^{1}$. Actually, it is not even Lipschitz, but almost Lipschitz. In the sense of definition (3) the set $M$ consists of the interior points of the side and the bases. The complement $M_{0}$ of $M$ is the union of the boundaries of the bases. Here the extension (3) of the Divergence Theorem applies as the exceptional set $M_{0}$ is small enough, i.e. its $3$-dimensional Minkowski content is zero.
(9) Lemma. Let $\tau:\mathbb{R}^{3}\to\mathbb{R}$ and $S:=\{(\tau(x),x):x\in\mathbb{R}^{3}\}$. Then $S$ is a spacelike Cauchy surface if $|\tau(x)-\tau(y)|<|x-y|$ for $x\neq y$ and $\limsup_{|x|\to\infty}|\tau(x)|/|x|<1$ or if a fortiori $\tau$ is $L$-Lipschitz with $L<1$, in which case $S$ is called $L$-achronal.
@@
(10) Theorem. Let the real $C^{1}$ vector field $\mathfrak{j}$ on $\mathbb{R}^{4}$ be bounded and satisfy the continuity equation (4.1). Suppose that $\mathfrak{j}$ is zero or causal future-directed, i.e.
@@
       (b) The claim is that $\mathfrak{h}$ is locally Lipschitz. By [13, Proposition 37], $\Phi:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}^{4}$, $\Phi(s,x):=\gamma_{x}(s)$ is a diffeomorphism. Write $\Phi=:(\varphi,\phi)$.
       From the implicit function theorem [20] applied to the function $f:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}$, $f:=\varphi-\tau\circ\phi$, it follows that $\sigma:\mathbb{R}^{3}\to\mathbb{R}$ is continuous and hence that $\mathfrak{h}$ is continuous.
       Indeed, since $f(s,x)=0$ is uniquely solved by $s=\sigma(x)$ for every $x\in\mathbb{R}^{3}$ it remains to verify that $f(\cdot,x)$ is injective for every $x\in\mathbb{R}^{3}$. Let $f(s,x)=f(s^{\prime},x)$. This implies $|\varphi(s,x)-\varphi(s^{\prime},x)|\leq|\phi(s,x)-\phi(s^{\prime},x)|$. Since $\gamma_{x}$ is timelike, $s=s^{\prime}$ follows.
       Now we show that $\mathfrak{h}$ is even locally Lipschitz. Put $\mathfrak{G}:=(\varphi,\operatorname{id}_{\mathbb{R}^{3}})$. As $\partial_{s}\varphi(s,x)=v_{0}(\Phi(s,x))>0$ one infers that $\mathfrak{G}$ is an diffeomorphisms. It follows $g(h_{0}(x),x)=h(x)$ for $g:=\phi\circ\mathfrak{G}^{-1}$.
       Fix $x,y\in\mathbb{R}^{3}$. Define $\beta:[0,1]\to\mathbb{R}^{3}$, $\beta(u):=g(\mathfrak{z}(u))$ with $\mathfrak{z}(u):=\big{(}(h_{0}(y),y)-(h_{0}(x),x)\big{)}u+(h_{0}(x),x)$. Note $\beta(1)=h(y)$, $\beta(0)=h(x)$ and remind $\mathfrak{h}(\mathbb{R}^{3})=S$.
@@
       Let $D_{\text{time}}g$ denote the $\mathbb{R}^{3}$-valued function obtained by differentiating each component of $g$ with respect to its first variable, and let $D_{\text{space}}g$ denote the $3\times 3$ matrix-valued function given by the Jacobian of $g$ with respect to its last three variables. Then $\beta^{\prime}(u)=D_{time}g(\mathfrak{z}(u))\big{(}(h_{0}(y)-h_{0}(x)\big{)}+D_{space}g(\mathfrak{z}(u))\big{(}y-x\big{)}$, whence
$|h(y)-h(x)|\leq|D_{time}g(\mathfrak{z}(u_{0})|\,|h(y)-h(x)|+||D_{space}g(\mathfrak{z}(u_{0}))||\,|y-x|$
@@
       Let $K\subset\mathbb{R}^{3}$ be a compact ball. The image of $K$ under the continuous map $(h_{0},\operatorname{id}_{\mathbb{R}^{3}})$ is contained in a compact ball $D\subset\mathbb{R}^{4}$, which is convex. Hence for $x,y\in K$ one has $\mathfrak{z}(u)\in D$ and therefore in conclusion $|h(y)-h(x)|\leq C_{D}|y-x|$.
       (c) Let $0<R<\infty$. Put $A_{R}:=\{x\in\mathbb{R}^{3}:|x|=R\}$ and $B_{R}:=\{x\in\mathbb{R}^{3}:|x|<R\}$. Suppose that $\sigma(x)>0$ for $|x|<R$. Then
$E_{R}:=\bigcup_{x\in B_{R}}\{\gamma_{x}(s):0<s<\sigma(x)\}$
is open in $\mathbb{R}^{4}$ with $\partial E_{R}=M\cup M_{0}$, where $M:=\{0\}\times B_{R}\cup L_{R}\cup\mathfrak{h}(B_{R})$ with $L_{R}:=\bigcup_{x\in A_{R}}\{\gamma_{x}(s):0<s<\sigma(x)\}$, and $M_{0}:=\{0\}\times A_{R}\cup\mathfrak{h}(A_{R})$.
       Now the proof of the basic idea in (d)-(f) follows. One verifies that the divergence theorem (3) applies to $n=4$, $E=E_{R}$ showing $\mathcal{H}^{3}(\partial E_{R})<\infty$ by (d), (e), and $\mathcal{M}^{3}(M_{0})=0$ by (f), (3). It is decisive that the integral over $L_{R}$ is zero as $L_{R}$ is constituted of integral curves.
       (d) Let $C\subset\mathbb{R}^{3}$ be bounded Borel. Then $\mathcal{H}^{3}(\{0\}\times C)<\infty$, $\mathcal{H}^{3}(\operatorname{graph}\tau|_{C})<\infty$ by [21, Theorem 8.1].
@@
       Recall the diffeomorphism $\Phi:\mathbb{R}\times\mathbb{R}^{3}\to\mathbb{R}^{4}$. Hence $f:=\mathbb{R}\times]-R/2,R/2[\times]-R/2,R/2[\to\mathbb{R}^{4}$, $f(s,x_{1},x_{2}):=\Phi\big{(}s;x_{1},x_{2},\sqrt{R^{2}-x_{1}^{2}-x_{2}^{2}}\,\big{)}$ is injective, $C^{1}$ whence Lipschitz on compact sets, and $f([-S,S]\times[-R/3,R/3]\times[-R/3,R/3])=D^{\prime}$. The claim holds by [21, Theorem 8.1].
       (f) Let $C\subset\mathbb{R}^{3}$ be a compact Lebesgue null set. Then $\mathcal{H}^{3}(\{0\}\times C)=0$ and $\mathcal{H}^{3}(\mathfrak{h}(C))=0$ hold by [21, Theorem 8.1]. Indeed, the former is obvious, the latter holds true as $\mathfrak{h}$ is Lipschitz on $C$ being compact.
       (g) Now let $\eta\in C_{c}^{1}(\mathbb{R}^{4})$ with $0\leq\eta\leq 1$ and $\eta|_{\overline{E}_{R}}=1$. Put $\phi:=\eta\mathfrak{v}$. Then (3) yields $0=\int_{M}\mathfrak{v}\,\nu_{E_{R}}\operatorname{d}\mathcal{H}^{3}$ as $\mathfrak{v}$ satisfies the continuity equation. Moreover the integration over $L_{R}$ yields 00 as the integrand is 00. So using [21, Theorem 9.1] it follows $\int_{\varpi(\mathfrak{h}(B_{R}))}\big{(}v_{0}(\tau(x),x)-v(\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}x^{3}=\int_{B_{R}}v_{0}(0,x)\operatorname{d}x^{3}$.
       (h) Finally, as explained in [13, Proposition 37], the condition $\sigma(x)>0$ for $|x|<R$ in (c) can be removed and the limit $R\to\infty$ carried out. One ends up with (4.2).∎
Under the assumptions made on the vector field $\mathfrak{j}$, the result in (4) states that the future-directed flux passing through a spacelike Cauchy surface is the same for all these surfaces. Under the additional assumption (4)(b) on $\mathfrak{j}$ the result in (4) is extended to all maximal achronal sets containing the origin simply by flatten them becoming $\gamma$-achronal for $\gamma<1$.
(11) Lemma. Let the real $C^{1}$ vector field $\mathfrak{j}$ be bounded and satisfy the continuity equation (4.1). Suppose that
@@
    $j_{0}(\mathfrak{x})\geq|j(\mathfrak{x})|$ for $\mathfrak{x}\in\mathbb{R}^{4}$
@@
    $j_{0}(\mathfrak{x})\leq C(1+|x|)^{-N}$ for $|x|\geq|x_{0}|$ with some constants $N>3$ and $C<\infty$
Then for every maximal achronal set $\Lambda$ with $0\in\Lambda$
@@
holds. Here $\operatorname{graph}\tau=\Lambda$ with $\tau(0)=0$. The integrands are nonnegative and the integrals are finite.
Proof. Due to (a) and $|\nabla\tau|\leq 1$ the integrands are nonnegative. Note $|\tau(x)|=|\tau(x)-\tau(0)|\leq|x-0|=|x|$. Hence, by (b), $\big{|}j_{0}(\tau(x),x)-j(\tau(x),x)\nabla\tau(x)\big{|}\leq 2j_{0}({\tau(x)},x)\leq 2C_{N}(1+|x|)^{-N}$ with $N>3$, whence the integrals are finite.
       Let $10<\gamma<$ 1. Then $\mathfrak{j}(\gamma x_{0},x)\to\mathfrak{j}(x_{0},x)$ for $\gamma\to 1$ by continuity. Note that $\gamma\tau$ is $\gamma$-Lipschitz. Hence by (4) the corresponding maximal achronal set is a spacelike Cauchy surface. Therefore by (4) equation (4.3) holds for $\gamma\tau$ in place of $\tau$. Note that still $|\gamma\tau(x)|\leq|x|$. Thus, by (b), the map $\mathbb{R}^{3}\ni x\mapsto 2C_{N}(1+|x|)^{-N}$ is an integrable majorant uniform with respect to $\gamma$, whence the claim by dominated convergence.∎
@@ 5 Covariant achronal localization out of covariant conserved current
       There exist AL with the localization operators $T(\Delta)$ being orthogonal projections [8, (22) Theorem]. In this case the localization operators commute. However, a quantum mechanical system localized by a projection valued AL necessarily does not have a semi-bounded energy operator. This no-go result following from Hegerfeldt’s well-known theorem [17] [^3] See also [18] and D.Buchholz and J.Yngvason’s comment also on Hegerfeld’s achievements [2]. regards a first type of Einstein causality requirement whose modern generalized reformulation [6, 8] we shall present below.
       The notion of localization of the above definition seems appropriate to describe measurement processes where a quantum system is absorbed by the apparatus and no further localization measurments can be performed on it. This is because, in case of subsequent measurements a second type of Einstein’s causality requirement essentially regarding the no-signaling condition has to be considered: The statistics of the measurement results $\Delta$ must be the same, regardless of whether a non-selective measurement was made in $\Delta^{\prime}$ or not.
To comply with this requirement, under some popular assumptions about the post measurement state, $T(\Delta)$ and $T(\Delta^{\prime})$ should commute if $\Delta$ and $\Delta^{\prime}$ cannot be joined by causal curves. [^4] If we explicitly assume that $T(\Delta)$ and $T(\Delta^{\prime})$ are projectors and suppose that the post-measurement state of a non-selective measurement of $T(\Delta^{\prime})$ and $I-T(\Delta^{\prime})$ is obtained with the Lüders-von Neumann projection postulate for every initial state, the no-signaling requirement easily implies that $T(\Delta)$ and $T(\Delta^{\prime})$ commute. If $T(\Delta)$ and $T(\Delta^{\prime})$ are effects, the no-signaling condition is guaranteed if, more weakly, the respective Kraus operator commute. However, if we also assume, as is common in local quantum physics, that all relevant operators (including Kraus ones) are organized in local von Neumann algebras (here associated to the causal completions of $\Delta$ and $\Delta^{\prime}$), the Kraus operator of $T(\Delta)$ must also commute with the adjoint of the Kraus operator of $T(\Delta^{\prime})$. In summary, $T(\Delta)$ and $T(\Delta^{\prime})$ commute. This requirement cannot be fulfilled as a consequence of Malament’s theorem [22] and its modern re-formulations due to Halvorson and Clifton particular [16]. A quick review on these issues related to various facets of Einstein’s causality and localization appears in the introduction of [23].
@@
In [13], condition CC was proved true in a special case where the sets $\Delta$ belong to smooth Cauchy surfaces and in [8], the final result has been established that an AL satisfies CC in full. As argued in [8], CC even necessitate achronal localization. In fact spacelike localization is not sufficient since CC induces the localization in achronal hyperplanes. This fact is reported in [8] and studied in detail in [9]. An important result is that the concepts of achronal localization AL and representation of the causal logic RCL are equivalent (cf. (7)) and that just the existence of a maximal spacelike set, which is not maximal achronal because of a missing piece of an achronal not spacelike hyperplane (see $P$ in [8, (8) Example],) is the reason behind the missing orthomodularity of the lattice of the causally complete sets generated by the spacelike relation (cf. [6, sec. 11.3]).
       Hence we consider the study and the explicit construction of the achronal localization of the massive scalar boson to be of utmost relevance.
       It is the very principle of causality which lets one think of the probability of localization as a conserved quantity reigned by an associated density current $\mathfrak{J}$. Indeed, by (5) a covariant AL can be constructed by means of a covariant conserved current assuming that on the Euclidean space its zeroth component describes the density of the probability of localization, namely explicitly $J_{0}(\phi;0,x)\geq 0$, $x\in\mathbb{R}^{3}$, and
@@
(13) Definition. Let $W$ be a unitary representation of $\tilde{\mathcal{P}}$ and let $\mathcal{D}$ be a $W$-invariant dense subspace of $\mathcal{H}$. Let $\mathfrak{J}=(J_{0},J)$ be a map from $\mathcal{D}\times\mathbb{R}^{4}$ to $\mathbb{R}^{4}$ such that $\mathfrak{J}(\phi,\cdot)$ is a bounded $C^{1}$ vector field. (i) $\mathfrak{J}$ is a conserved current if the latter satisfies the continuity equation. (ii) $\mathfrak{J}$ is a (Poincaré) covariant current if $\mathfrak{J}\big{(}W(g)\phi,\mathfrak{x}\big{)}=A\cdot\mathfrak{J}\big{(}\phi,g^{-1}\cdot\mathfrak{x}\big{)}$, $g=(\mathfrak{a},A)\in\tilde{\mathcal{P}}$ holds.
@@
(18) Lemma. Let $\mathfrak{J}$ be conserved and covariant. Let $\phi\in\mathcal{D}$. Suppose (4)(a),(b) for $\mathfrak{J}(\phi,\cdot)$ and suppose $\int J_{0}(\phi;0,x)\operatorname{d}^{3}x=||\phi||^{2}$. Then
$||\phi||^{2}=\int\big{(}J_{0}(\phi;\tau(x),x)-J(\phi;\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}^{3}x$
holds for every maximal achronal set with corresponding $1$-Lipschitz function $\tau$.
Proof. By covariance of $\mathfrak{J}$ regarding time translations it follows that $\mathfrak{J}(\phi;\tau(x),x)=\mathfrak{J}(W(-\tau(0))\phi;\tau(x)-\tau(0),x)=\mathfrak{J}(\phi^{\prime};\tau^{\prime}(x),x)$ for $\phi^{\prime}:=W(-\tau(0))\phi$, $\tau^{\prime}:=\tau-\tau(0)$. Hence the right side of (4.3) reads $\int J_{0}(\phi^{\prime};0,x)\operatorname{d}^{3}x=||\phi^{\prime}||^{2}=||\phi||^{2}$.∎
@@
(19) Theorem. Let the real bounded $C^{1}$ current $\mathfrak{J}$ be conserved covariant. Let $\phi\in\mathcal{D}$. Let $J_{0}$ satisfy
@@
    $J_{0}(\phi,\mathfrak{x})\leq C(1+|x|)^{-N}$ for $|x|\geq|x_{0}|$ with some constants $N>3$ and $C<\infty$ depending on $\phi$
Then there is a unique AL $T$ satisfying for every achronal Borel set $\Delta$
@@
Uniqueness is obvious as $\mathcal{D}$ is dense. Let $\phi\in\mathcal{D}$.
       By (a) and (5), $J_{0}(\phi,\mathfrak{x})\geq|J(\phi,\mathfrak{x})|$.